# FastDraft: How to Train Your Draft

### v1

@@ Abstract
Speculative Decoding has gained popularity as an effective technique for accelerating the auto-regressive inference process of Large Language Models (LLMs). However, Speculative Decoding entirely relies on the availability of efficient draft models, which are often lacking for many existing language models due to a stringent constraint of vocabulary incompatibility. In this work we introduce FastDraft, a novel and efficient approachfor pre-training and aligning a draft model to any large language model by incorporating efficient pre-training, followed by fine-tuning over synthetic datasets generated by the target model. We demonstrate FastDraft by training two highly parameter efficient drafts for the popular Phi-3-mini and Llama-3.1-8B models. Using FastDraft, we were able to produce a draft with approximately 10 billion tokens on a single server with 8 Intel® Gaudi® 2 accelerators in under 24 hours. Our results show that the draft model achieves impressive results in key metrics of acceptance rate, block efficiency and up to 3x memory bound speed up when evaluated on code completion and up to 2x in summarization, text completion and instruction tasks. We validate our theoretical findings through benchmarking on the latest Intel® Core Ultra, achieving a wall-clock time speedup of up to 2x, indicating a significant reduction in runtime. Due to its high quality, FastDraft unlocks large language models inference on AI-PC and other edge-devices.
@@ 1 Introduction
The advent of Transformer architectures has fundamentally reshaped the field of natural language processing (NLP). In recent years, Transformer-based models have achieved remarkable success across a broad spectrum of natural language understanding and generation tasks [Achiam et al., 2023, Team et al., 2023, Abdin et al., 2024]. Their exceptional performance, particularly in large language models (LLMs), has made them highly desirable for deployment in numerous applications, ranging from conversational systems to content generation and beyond. Despite their outstanding performance, LLMs suffer from slow inference speeds due to substantial memory bandwidth requirements and the sequential nature of auto-regressive generation (ARG). The introduction of Speculative Decoding (SD) [Leviathan et al., 2023] offers a promising solution for accelerating ARG without sacrificing generation quality, making it a compelling approach for improving LLM inference efficiency. SD utilizes a draft language model (LM) to generate a sequence of tokens auto-regressively, while the target model validates the batched tokens in parallel. In certain applications, SD can achieve a 2-3x speedup in LLM inference without compromising the generation quality of the target model. Achieving significant speedup with SD requires a high-quality draft model that is both efficient and well-aligned with the target. To date, such draft models remain scarce, even for widely used open-source LLMs [Dubey et al., 2024, Abdin et al., 2024] due to vocabulary incompatibility. To address this limitation, we propose FastDraft, a method for producing hardware-efficient draft models that are orders of magnitude smaller than their corresponding target models.
@@
While extensive research has focused on training and data generation for high-quality LLMs [Kaplan et al., 2020, Longpre et al., 2023], these frameworks are not necessarily applicable to training draft models for SD. LLMs are typically trained to generate helpful, high-quality responses, whereas draft models should be trained to generate sequences that are likely to be accepted by the target model. We explore previously unexamined aspects of draft model training for SD. Our contributions include: (1) introducing a method for producing quality and highly efficient draft models with low resource requirements for any given target LLM and demonstrate it by training and benchmarking a draft for Phi-3-mini, (2) conducting extensive ablation studies on pre-training data size, pre-training for both code and natural language and target-draft alignment via knowledge distillation (KD) and (3) demonstrating the scalability of FastDraft by training a draft model for Llama-3.1-8B-Instruct, achieving performance improvements comparable to those attained for Phi-3-mini. We demonstrate significant improvements in key metrics using FastDraft, leading to theoretical speedups of up to 3x as measured by the Memory Bound Speedup (MBSU) metric. We further validate our theoretical findings by evaluating the wall-clock time speedup achieved with our approach on the latest Intel® Core Ultra. Our results demonstrate an average 1.5x speedup for natural language tasks and an average 2x speedup for code completion tasks. According to our findings, the small size of the draft model and the limited amount of data required to produce a high-quality draft enabled us to successfully train and align a draft model to Phi-3-mini, end-to-end, in under 24 hours using a single node with 8 Intel® Gaudi® 2 accelerators.
@@ 2 Related work
The widespread adoption of LLMs in cloud and edge devices has driven a significant body of research focused on developing alternative strategies for ARG to address the slow performance of LLM inference [Santilli et al., 2023, Ghazvininejad et al., 2019, Stern et al., 2018]. However, many of these approaches compromise generation quality or require additional training data and architectural modifications. The introduction of speculative decoding as a lossless solution for accelerating LLM inference [Leviathan et al., 2023] has inspired a new wave of follow-up research. Some studies propose using plug-in prediction heads as a drafting mechanism [Zhang et al., 2024, Cai et al., 2024], while others focus on improving the serving latency of stand-alone draft models [Sun et al., 2024, Chen et al., 2023, Miao et al., 2023]. In contrast, our work focuses on directly enhancing stand-alone draft model’s capabilities through pre-training and fine-tuning. Other works apply KD to draft models to improve alignment with the target model. These studies explore various divergence functions for the KD algorithm, rather than relying solely on the commonly used Kullback–Leibler Divergence (KLD). For instance, Zhou et al. [2023] proposed Total Variation Distance (TVD) based on Corollary 3.6 in [Leviathan et al., 2023], which posits that minimizing TVD maximizes the token-level acceptance rate. Goel et al. [2024] further developed this approach with TVD++, fine-tuning their model, pre-trained on 600 billion tokens. Another study, Yan et al. [2024], focuses on enhancing the hardware efficiency of draft models by extensively analyzing the trade-off between time latency and acceptance rates, rather than relying solely on the latter. While their work emphasizes optimizing the draft architecture, our approach fixes the architecture and introduces efficient approach for pre-training and aligning a draft model. Li et al. [2024] proposes training a compact draft model based on the target model, using a relatively limited dataset. The primary aim of the paper is to develop a draft architecture that effectively utilizes the target model’s hidden representations and weights. In contrast, our paper centers on training and aligning any draft architecture that shares only the vocabulary with the target model.
@@ 3 Speculative decoding
SD is a lossless decoding paradigm introduced by Leviathan et al. [2023] for accelerating ARG with LLMs. It is inspired by speculative execution [Hennessy and Patterson, 2017] and aims to mitigate the inherent latency bottleneck caused by the sequential nature of ARG [Pope et al., 2023]. SD employs a draft LM to generate a block of $\gamma$ candidate tokens. The LLM, referred to as the target model, then processes these candidate tokens in parallel. The algorithm examines each token’s probability distribution, calculated by both the target and draft models, to determine whether the token should be accepted or rejected. As a result, any LM can function as a draft model, provided it shares the same vocabulary as the target model. However, since an LM’s vocabulary is fixed during pre-training, leveraging existing models as draft models is only feasible if they were pre-trained on the same vocabulary. A key metric for benchmarking ARG inference performance is time per output token (TPOT), which represents the expected latency for generating a single output token, excluding pre-filling latency.1 1 1 Pre-filling refers to the action of populating the key-value cache with the input tokens’ information. In SD, TPOT is a function of the draft model latency $l_{D}$, the speculation block size $\gamma$, the corresponding expected block efficiency $\tau^{\gamma}$2 2 2 The expected number of tokens accepted within a speculation block., and the target model latency $l_{T}^{\gamma}$. The equation is expressed as:
@@ 4.1 Draft architecture & pre-training
The draft architecture imposes only one strict requirement, it must produce a probability distribution over the target’s vocabulary. Beyond this, the design is flexible. Nevertheless, certain factors should be considered when selecting the draft’s architecture, with latency being the primary concern, as discussed in Section 3. Then, the chosen draft is trained over a pre-training dataset of natural language for language modeling with the objective of predicting the next token. A common application of LLMs is code completion. While the pre-training dataset may include some code, unless the dataset is specifically focused on code, the draft’s performance on tasks requiring an understanding of code is often suboptimal. To address this, Aryabumi et al. [2024] proposed continued pre-training (CP), where training begins with a pre-trained model and is extended using a combination of code and natural language data. FastDraft adopts this approach for producing drafts for code completion tasks.
@@ 4.2 Target-draft alignment
One of the primary objectives when designing a draft model is to maximize the acceptance rate with the target model, as a higher acceptance rate directly leads to greater speedup in SD. To closely mimic the target model’s behavior in real-world scenarios, we investigate two knowledge distillation strategies that expose the draft model to data samples that closely reflect those generated by the target model.
@@
Kim and Rush [2016] proposed performing knowledge distillation (KD) by training the student model on sequences generated by the teacher model. This approach is widely used in large language model (LLM) training to improve model quality [Taori et al., 2023, Abdin et al., 2024]. To align the pre-trained draft model with the teacher model, we employ sequence-level KD by fine-tuning the draft model on a synthetic dataset generated by the target. In contrast to the sequence-level KD approach proposed by Kim and Rush [2016], we differentiate between training on teacher-generated data and training with knowledge distillation using the teacher’s logits. Accordingly, in this work, sequence-level KD refers specifically to training on generated tokens with cross-entropy loss, without incorporating the teacher’s soft targets (logits).
@@
This strategy aligns with the traditional KD method presented in [Hinton, 2015], which involves calculating the divergence of the token level probability distribution over the vocabulary between the teacher and the draft. Since computing the KD loss depends on the teacher’s logits, the teacher is required to perform inference at every training step of the draft, making this process both computationally intensive and time-consuming. An alternative approach is to precompute the teacher’s logits beforehand and embed them into the dataset. However, the memory demands of such a dataset can quickly exceed several terabytes, as the number of logits per token equals the vocabulary size3 3 3 The vocabulary size usually exceeds 30,000 tokens posing a substantial memory overhead. To mitigate this issue, only a small subset of the most significant logits per token is extracted, significantly reducing the memory requirements. This optimization results in a 6x-9x reduction in fine-tuning time without noticeable degradation in model quality.
@@ 4.3.1 Metrics
Given the block efficiency $\tau^{\gamma}$, the expected speedup achieved by applying Speculative Decoding (SD) is expressed as $\frac{\tau^{\gamma}}{c\gamma+1}$, where $c$ represents the ratio of latencies between the draft and target models (see Equation 2 or [Leviathan et al., 2023]). Since this metric is hardware-dependent, it is preferable to use a hardware-agnostic measure. Considering that the expected speedup occurs in a memory-bound regime, we utilize the Memory-Bound Speedup (MBSU). We define $\hat{c}$ as the ratio of parameter counts between the draft and target models.
@@ 4.3.2 Benchmarks
The most commonly used benchmarks for assessing the quality of draft models, typically open-ended generation and summarization tasks, include the XSum [Narayan et al., 2018], CNN/DailyMail [Hermann et al., 2015], and HumanEval [Chen et al., 2021] datasets. In addition to CNN/DailyMail and HumanEval, we include in our evaluation the TinyStories [Eldan et al., 2023] and Dolly [Conover et al., 2023] datasets, providing a more comprehensive assessment of model performance across diverse tasks. When utilizing the TinyStories dataset, we generate text by starting from a random position within each sample, using the preceding tokens as the input context.
Table 1: Pre-training data size effect on acceptance rate and perplexity. Results for block size $\gamma=3$ and multinomial sampling with temperature $T=0.6$
Draft size Data size Perplexity CNN-DailyMail TinyStories Dolly
@@
5 Experiments
5.1 Experimental setup
We demonstrate FastDraft by training a draft model for Phi-3-mini-4k-Instruct  Abdin et al. [2024] target model. Phi-3-mini, a 3.8 billion parameters LLM, was selected as our case study due to its outstanding performance across multiple open-source LLM benchmarks, while also being capable of running locally on edge devices, such as personal computers and smartphones.
Draft architecture
Pre-training datasets
For our natural language dataset, we use a 10 billion tokens (BT) sample from the FineWeb dataset [Penedo et al., 2024]. FineWeb is a widely used open-source, de-duplicated, and quality-filtered dataset, comprising 15 trillion tokens derived from 96 Common Crawl snapshots [Crawl, ]. It has been demonstrated to yield better-performing LLMs compared to other open-source pre-training datasets such as C4, RedPajama, and The Pile (Section 3.7 in Penedo et al. [2024]). For our code dataset, we use a 10BT sample from The Stack v2 smol dataset [Lozhkov et al., 2024]. This variant of The Stack v2 consists of 17 commonly used programming languages, as well as a substantial collection of documentation languages, configuration languages, and configuration files. An overview of the 10BT dataset composition is provided in Appendix A.2.
Synthetic alignment dataset
To perform sequence-level KD as discussed in Section 4.2, we produce an instruction fine-tuning collection using the seed instructions from several open instruction datasets: Alpaca [Taori et al., 2023], OIG-small-chip2 [Huu et al., 2023] and Evol-Instruct [Luo et al., 2023]. We collect response sequences generated by Phi-3-mini with greedy sampling along with multinomial sampling with temperature in $\{0.6,0.8,1.0\}$ to improve the diversity of the generated sequences. Additionally, we adopt the approach proposed by Xu et al. [2024], directly soliciting instructions from the target model. These instructions are then used to generate the corresponding responses in the same manner as previously described.
Figure 1: MBSU of 50M and 120M pre-trained drafts on 5BT FineWeb sample.
5.2 Pre-training dataset size
We study the impact of the pre-training dataset size on the performance of the draft model, aiming to optimize runtime and resource usage and prevent diminishing returns. Consequently, we uniformly sample subsets of sizes $\{0.1,0.5,1,2,5,10\}$BT and pre-train the 50M and 120M draft configurations on these subsets. We report the AR results for the $\{2,5,10\}$BT subsets with block size $\gamma=3$ and multinomial sampling in addition to perplexity results measured on Wikitext2 [Salesforce, ] in Table 1. The full results are reported in Table 7 in the Appendix.
As anticipated, perplexity values decrease as the models are exposed to more training data, though the rate of improvement slows down. However, when looking at the AR results, it is not the case. In CNN-DailyMail and Dolly AR either plateaus or decreases as the data size grows. In the case of TinyStories, AR increases with the amount of data which can be expected due to the nature of the benchmark of text completion versus instruction following and summarization. Overall, we obtain strong results across all dataset sizes, with the models trained on the 5BT dataset emerging as a promising middle-ground option. When comparing the 50M draft to the 120M draft, the 120M draft demonstrates superior performance in terms of AR. Additionally, while the 120M draft also outperforms the 50M model in MBSU, the margin of improvement is relatively small, see Figure 1. However, it is important to note that training the 120M draft requires twice the time compared to the 50M draft.
For the subsequent experiments, we selected the 50M draft model trained on 5BT tokens as our pre-trained draft. Unless otherwise specified, references to a pre-trained draft in this section pertain to this model.
Figure 2: Block efficiency results for continued pre-training with code on tasks: CNN-DailyMail, TinyStories, Dolly and HumanEval using greedy decoding with block size $\gamma=3$

### v2

@@ Abstract
Speculative Decoding has gained popularity as an effective technique for accelerating the auto-regressive inference process of Large Language Models. However, Speculative Decoding entirely relies on the availability of efficient draft models, which are often lacking for many existing language models due to a stringent constraint of vocabulary compatibility. In this work we introduce FastDraft, a novel and efficient approach for pre-training and aligning a draft model to any large language model by incorporating efficient pre-training, followed by fine-tuning over synthetic datasets generated by the target model. We demonstrate FastDraft by training two highly parameter efficient drafts for the popular Phi-3-mini and Llama-3.1-8B models. Using FastDraft, we were able to produce a draft model with approximately 10 billion tokens on a single server with 8 Intel® Gaudi® 2 accelerators in under 24 hours. Our results show that the draft model achieves impressive results in key metrics of acceptance rate, block efficiency and up to 3x memory bound speed up when evaluated on code completion and up to 2x in summarization, text completion and instruction tasks. We validate our theoretical findings through benchmarking on the latest Intel® Core Ultra, achieving a wall-clock time speedup of up to 2x, indicating a significant reduction in runtime. Due to its high quality, FastDraft unlocks large language models inference on AI-PC and other edge-devices.
FastDraft: How to Train Your Draft
Ofir Zafrir*, Igor Margulis*, Dorin Shteyman*, Shira Guskin, Guy Boudoukh Intel Labs {ofir.zafrir,igor.margulis,dorin.shteyman}@intel.com {shira.guskin,guy.boudoukh}@intel.com
@@ 1 Introduction
The advent of Transformer architectures has fundamentally reshaped the field of natural language processing (NLP). In recent years, Transformer-based models have achieved remarkable success across a broad spectrum of natural language understanding and generation tasks (Achiam et al., 2023; Team et al., 2023). Their exceptional performance, particularly in large language models (LLMs), has made them highly desirable for deployment in numerous applications, ranging from conversational systems to content generation and beyond. Despite their outstanding performance, LLMs suffer from slow inference speed due to substantial memory bandwidth requirements and the sequential nature of auto-regressive generation (ARG). The introduction of Speculative Decoding (SD) (Leviathan et al., 2023) offers a promising solution for accelerating ARG without sacrificing generation quality, making it a compelling approach for improving LLM inference efficiency. SD utilizes a draft language model (LM) to generate a sequence of tokens auto-regressively, while the target model validates the batched tokens in parallel. In certain applications, SD can achieve a 2-3x speedup in LLM inference without compromising the generation quality of the target model. Achieving significant speedup with SD requires a high-quality draft model that is both efficient and well-aligned with the target. To date, such draft models remain scarce, even for widely used open-source LLMs (Dubey et al., 2024; Abdin et al., 2024) due to vocabulary incompatibility.
To address this limitation, we propose FastDraft, a method for producing hardware (HW) efficient draft models that are orders of magnitude smaller than their corresponding target models.
@@
While extensive research has focused on training and data generation for high-quality LLMs (Kaplan et al., 2020; Longpre et al., 2023), these frameworks are not necessarily applicable to training draft models for SD. LLMs are typically trained to generate helpful, high-quality responses, whereas draft models should be trained to generate sequences that are likely to be accepted by the target model. We explore previously unexamined aspects of draft model training for SD. Our contributions include: (1) introducing a method for producing quality and highly efficient draft models with low resource requirements for any given target LLM and demonstrate it by training and benchmarking a draft for Phi-3-mini, (2) conducting extensive ablation studies on pre-training data size, pre-training for both code and natural language, target-draft alignment via knowledge distillation (KD) and HW-aware draft architectures, and (3) demonstrating the scalability of FastDraft by training a draft model for Llama-3.1-8B-Instruct, achieving performance improvements comparable to those attained for Phi-3-mini. We demonstrate significant improvements in key metrics using FastDraft, leading to theoretical speedups of up to 3x as measured by the Memory Bound Speedup (MBSU) metric. We further validate our theoretical findings by evaluating the wall-clock time speedup achieved with our approach on the latest Intel® Core Ultra. Our results demonstrate an average 1.5x speedup for natural language tasks and an average 2x speedup for code completion tasks. According to our findings, the small size of the draft model and the limited amount of data required to produce a high-quality draft enabled us to successfully train and align a draft model to Phi-3-mini, end-to-end, in under 24 hours using a single node with 8 Intel® Gaudi® 2 accelerators.
@@ 2 Related work
The widespread adoption of LLMs in cloud and edge devices has driven a significant body of research focused on developing alternative strategies for ARG to address the slow performance of LLM inference (Santilli et al., 2023; Ghazvininejad et al., 2019; Stern et al., 2018). However, many of these approaches compromise generation quality or require additional training data and architectural modifications. The introduction of speculative decoding as a lossless solution for accelerating LLM inference (Leviathan et al., 2023) has inspired a new wave of follow-up research. Some studies propose using plug-in prediction heads as a drafting mechanism (Zhang et al., 2024; Cai et al., 2024), while others focus on improving the serving latency of stand-alone draft models (Sun et al., 2024; Chen et al., 2023; Miao et al., 2023). In contrast, our work focuses on directly enhancing stand-alone draft model’s capabilities through pre-training and fine-tuning. Other works apply KD to draft models to improve alignment with the target model. These studies explore various divergence functions for the KD algorithm, rather than relying solely on the commonly used Kullback–Leibler Divergence (KLD). For instance, Zhou et al. (2023) proposed Total Variation Distance (TVD) based on Corollary 3.6 in (Leviathan et al., 2023), which posits that minimizing TVD maximizes the token-level acceptance rate. Goel et al. (2024) further developed this approach with TVD++, fine-tuning their model, pre-trained on 600 billion tokens. Another study, Yan et al. (2024), focuses on enhancing the HW efficiency of draft models by extensively analyzing the trade-off between time latency and acceptance rates, rather than relying solely on the latter. While their work emphasizes optimizing the draft architecture, our approach fixes the architecture and introduces efficient approach for pre-training and aligning a draft model. Li et al. (2024) proposes training a compact draft model based on the target model, using a relatively limited dataset. The primary aim of the paper is to develop a draft architecture that effectively utilizes the target model’s hidden representations and weights. In contrast, our paper centers on training and aligning any draft architecture that shares only the vocabulary with the target model.
@@ 3 Speculative decoding
SD is a lossless decoding paradigm introduced by Leviathan et al. (2023) for accelerating ARG with LLMs. It is inspired by speculative execution (Hennessy and Patterson, 2017) and aims to mitigate the inherent latency bottleneck caused by the sequential nature of ARG (Pope et al., 2023). SD employs a draft LM to generate a block of $\gamma$ candidate tokens. The LLM, referred to as the target model, then processes these candidate tokens in parallel. The algorithm examines each token’s probability distribution, calculated by both the target and draft models, to determine whether the token should be accepted or rejected. As a result, any LM can function as a draft model, provided it shares the same vocabulary as the target model. However, since an LM’s vocabulary is fixed during pre-training, leveraging existing models as draft models is only feasible if they were pre-trained on the same vocabulary. A key metric for benchmarking ARG inference performance is time per output token (TPOT), which represents the expected latency for generating a single output token, excluding pre-filling latency.1 1 1 Pre-filling refers to the action of populating the key-value cache with the input tokens’ information. In SD, TPOT is a function of the draft model latency $l_{D}$, the speculation block size $\gamma$, the corresponding expected block efficiency2 2 2 The expected number of tokens accepted within a speculation block. $\tau^{\gamma}$, and the target model latency $l_{T}^{\gamma}$. The equation is expressed as:
@@ 4.1 Draft architecture & pre-training
The draft architecture imposes only one strict requirement, it must produce a probability distribution over the target’s vocabulary. Beyond this, the design is flexible. Nevertheless, certain factors should be considered when selecting the draft’s architecture, with latency being the primary concern, as discussed in Section 3. Then, the chosen draft is trained over a pre-training dataset of natural language for language modeling with the objective of predicting the next token. A common application of LLMs is code completion. While the pre-training dataset may include some code, unless the dataset is specifically focused on code, the draft’s performance on tasks requiring an understanding of code is often suboptimal. To address this, Aryabumi et al. (2024) proposed continued pre-training (CP), where training begins with a pre-trained model and is extended using a combination of code and natural language data. FastDraft adopts this approach for producing drafts for code completion tasks.
@@ 4.2 Target-draft alignment
One of the primary objectives when designing a draft model is to maximize the acceptance rate with the target model, as a higher acceptance rate directly leads to greater speedup in SD. To closely mimic the target model’s behavior in real-world scenarios, we investigate two KD strategies that expose the draft model to data samples that closely reflect those generated by the target model.
@@
Kim and Rush (2016) proposed performing KD by training the student model on sequences generated by the teacher model. This approach is widely used in LLM training to improve model quality (Taori et al., 2023; Abdin et al., 2024). To align the pre-trained draft model with the teacher model, we employ sequence-level KD by fine-tuning the draft model on a synthetic dataset generated by the target. In contrast to the sequence-level KD approach proposed by Kim and Rush (2016), we differentiate between training on teacher-generated data and training with KD using the teacher’s logits. Accordingly, in this work, sequence-level KD refers specifically to training on generated tokens with cross-entropy loss, without incorporating the teacher’s soft targets (logits).
@@
This strategy aligns with the traditional KD method presented in (Hinton, 2015), which involves calculating the divergence of the token level probability distribution over the vocabulary between the teacher and the draft. Since computing the KD loss depends on the teacher’s logits, the teacher is required to perform inference at every training step of the draft, making this process both computationally intensive and time-consuming. An alternative approach is to precompute the teacher’s logits beforehand and embed them into the dataset. However, the memory demands of such a dataset can quickly exceed several terabytes, as the number of logits per token equals the vocabulary size3 3 3 The vocabulary size usually exceeds 30,000 tokens posing a substantial memory overhead. To mitigate this issue, only a small subset of the most significant logits per token is extracted, significantly reducing the memory requirements. This optimization results in a 6x-9x reduction in fine-tuning time without noticeable degradation in model quality.
@@ 4.3.1 Metrics
Given the block efficiency $\tau^{\gamma}$, the expected speedup achieved by applying Speculative Decoding (SD) is expressed as $\frac{\tau^{\gamma}}{c\gamma+1}$, where $c$ represents the ratio of latencies between the draft and target models. Since this metric is HW-dependent, it is preferable to use a HW-agnostic measure. Considering that the expected speedup occurs in a memory-bound regime, we utilize the Memory-Bound Speedup (MBSU). We define $\hat{c}$ as the ratio of parameter counts between the draft and target models.
@@ 4.3.2 Benchmarks
The most commonly used benchmarks for assessing the quality of draft models, typically open-ended generation and summarization tasks, include the XSum (Narayan et al., 2018), CNN-DailyMail (CNN-DM) (Hermann et al., 2015), and HumanEval (Chen et al., 2021) datasets. In addition to CNN-DM and HumanEval, we include in our evaluation the TinyStories (TS) (Eldan et al., 2023) and Dolly (Conover et al., 2023) datasets, providing a more comprehensive assessment of model performance across diverse tasks. When utilizing the TS dataset, we generate text by starting from a random position within each sample, using the preceding tokens as the input context.
5 Experiments
5.1 Experimental setup
We demonstrate FastDraft by training a draft model for Phi-3-mini-4k-Instruct  Abdin et al. (2024) target model. Phi-3-mini, a 3.8 billion parameters LLM, was selected as our case study due to its outstanding performance across multiple open-source LLM benchmarks, while also being capable of running locally on edge devices, such as personal computers and smartphones.
Draft architecture
Pre-training datasets
For our natural language dataset, we use a 10 billion tokens (BT) sample from the FineWeb dataset (Penedo et al., 2024). FineWeb is a widely used open-source, de-duplicated, and quality-filtered dataset, comprising 15 trillion tokens derived from 96 Common Crawl snapshots (Crawl, ). It has been demonstrated to yield better-performing LLMs compared to other open-source pre-training datasets such as C4, RedPajama, and The Pile (Section 3.7 in Penedo et al. (2024)). For our code dataset, we use a 10BT sample from The Stack v2 smol dataset (Lozhkov et al., 2024). This variant of The Stack v2 consists of 17 commonly used programming languages, as well as a substantial collection of documentation languages, configuration languages, and configuration files. An overview of the 10BT dataset composition is provided in Appendix A.2.
Synthetic alignment dataset
To perform sequence-level KD as discussed in Section 4.2, we produce an instruction fine-tuning collection using the seed instructions from several open instruction datasets: Alpaca (Taori et al., 2023), OIG-small-chip2 (Huu et al., 2023) and Evol-Instruct (Luo et al., 2023). We collect response sequences generated by Phi-3-mini with greedy sampling along with multinomial sampling with temperature in $\{0.6,0.8,1.0\}$ to improve the diversity of the generated sequences. Additionally, we adopt the approach proposed by Xu et al. (2024), directly soliciting instructions from the target model. These instructions are then used to generate the corresponding responses in the same manner as previously described.
5.2 Pre-training dataset size
We study the impact of the pre-training dataset size on the performance of the draft model, aiming to optimize runtime and resource usage and prevent diminishing returns. Consequently, we uniformly sample subsets of sizes $\{0.1,0.5,1,2,5,10\}$BT and pre-train the 50M and 120M draft configurations on these subsets. We report the AR results for the $\{2,5,10\}$BT subsets with block size $\gamma=3$ and multinomial sampling in addition to perplexity results measured on Wikitext2 (Salesforce, ) in Table 1. The full results are reported in Table 7 in the Appendix.
Draft Data size PPL CNN-DM TS Dolly
@@ 5.2 Pre-training dataset size
Table 1: Pre-training data size effect on AR and perplexity (PPL). Results for block size $\gamma=3$ and multinomial sampling with temperature $T=0.6$
As anticipated, perplexity values decrease as the models are exposed to more training data, though the rate of improvement slows down. However, when looking at the AR results, it is not the case. In CNN-DM and Dolly AR either plateaus or decreases as the data size grows. In the case of TS, AR increases with the amount of data which can be expected due to the nature of the benchmark of text completion versus instruction following and summarization. Overall, we obtain strong results across all dataset sizes, with the models trained on the 5BT dataset emerging as a promising middle-ground option. When comparing the 50M draft to the 120M draft, the 120M draft demonstrates superior performance in terms of AR. Additionally, while the 120M draft also outperforms the 50M model in MBSU, the margin of improvement is relatively small, see Figure 1. However, it is important to note that training the 120M draft requires twice the time compared to the 50M draft.
For the subsequent experiments, we selected the 50M draft model trained on 5BT tokens as our pre-trained draft, unless otherwise specified.
Figure 1: MBSU of 50M and 120M drafts pre-trained on 5BT FineWeb sample.
Figure 2: Block efficiency results for continued pre-training with code on tasks: CNN-DM, TS, Dolly and HumanEval using greedy decoding with block size $\gamma=3$